OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @mathmath333

OpenStudy (anonymous):

\[Given~\log _{3}m=2-\log _{3}n,\]express m in terms of n.

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\log _{3}m=2-\log _{3}n \hspace{.33em}\\~\\ \log _{3}m+\log _{3}n=2 \hspace{.33em}\\~\\ \log _{3}(m\times n)=2 \hspace{.33em}\\~\\ 3^2=m\cdot n \hspace{.33em}\\~\\ m=\dfrac{9}{n} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Beautiful :)

OpenStudy (anonymous):

Thnx @mathmath333 :)

OpenStudy (mathmath333):

yw